How many liters each of a 55% acid solution and a 80% acid solution must be used to produce 50 liters of a 60% acid solution? (Round to two decimal places if necessary.)

Answer by addingup(3677)

0.55x+0.8(50-x) = 0.6(50)
0.55x+40-0.8x = 30
-0.25x = -10 divide both sides by -0.25 and remember -/- = +
x = 40
You need 40 gallon of the 55% acid solution and 50-40 = 10 gallon of the 80% solution to get 50 gallon at 60%